# SpecGen: Neural Spectral BRDF Generation via Spectral-Spatial Tri-plane Aggregation

### v1

@@ Abstract
Synthesizing spectral images across different wavelengths is essential for photorealistic rendering. Unlike conventional spectral uplifting methods that convert RGB images into spectral ones, we introduce SpecGen, a novel method that generates spectral bidirectional reflectance distribution functions (BRDFs) from a single RGB image of a sphere. This enables spectral image rendering under arbitrary illuminations and shapes covered by the corresponding material. A key challenge in spectral BRDF generation is the scarcity of measured spectral BRDF data. To address this, we propose the Spectral-Spatial Tri-plane Aggregation (SSTA) network, which models reflectance responses across wavelengths and incident-outgoing directions, allowing the training strategy to leverage abundant RGB BRDF data to enhance spectral BRDF generation. Experiments show that our method accurately reconstructs spectral BRDFs from limited spectral data and surpasses state-of-the-art methods in hyperspectral image reconstruction, achieving an improvement of 8 dB in PSNR. Codes and data will be released upon acceptance.
@@
Figure 1: From a single RGB image of a sphere (left), our method recovers the spectral BRDF of the corresponding material, allowing the spectral image rendering of arbitrary illuminations and shapes covered by the material under different wavelengths (right). See Fig. 5 for more spectral renderings at more diverse wavelengths.
I Introduction
Real light is composed of a spectrum of wavelengths, and photorealistic image synthesis requires capturing the wavelength-dependent reflectance properties of materials under varying illumination and viewing conditions. Spectral bidirectional reflectance distribution functions (BRDFs) constitute a crucial ingredient in this process, enabling the synthesis of hyperspectral images for remote sensing [1, 2], material analysis [3], and virtual reality [4], as well as for classification tasks [5]. However, acquiring high-quality spectral BRDF data is challenging, as the measurement process requires scanning a four-dimensional domain at high resolution—a tedious and time-consuming task [6], leading existing spectral BRDF datasets to be scarce and limited in scale. To bypass the need for extensive spectral measurements, it is desired to develop a spectral BRDF generation method that can produce photorealistic spectral BRDFs from a single image input.
Existing methods have attempted to address BRDF generation, but they primarily focus on RGB BRDFs. For example, Gokbudak et al. [7] proposes a hypernetwork to reconstruct measured RGB BRDF from sparse samples. ControlMat [8] and MatFuse [9] leverage diffusion models to generate spatially-varying BRDFs, including albedo, roughness, metallic, and normal maps, from a single flash RGB photo. These methods, though effective in RGB BRDF generation, cannot be directly extended to spectral BRDFs due to differences in BRDF representation and the lack of spectral BRDF data. On the other hand, another line of research explores spectral uplifting [10, 11, 12, 13], which generates spectral images from RGB inputs. These methods mainly focus on producing spectral images, making extracting the underlying spectral BRDFs from the RGB inputs challenging. As a result, they cannot transfer the appearance of spectral material to different shapes and illuminations, a capability that is crucial for downstream image editing tasks. In summary, despite significant progress in spectral image generation and RGB BRDF generation, spectral BRDF generation remains an open problem.
@@ I Introduction
The key challenge of spectral BRDF generation lies in the scarcity of spectral BRDF data. Currently, only 51 isotropic spectral BRDFs are publicly available [6], which makes it difficult to train high-quality, generalizable models. In contrast, measured RGB BRDF datasets, such as MERL [14], are relatively easy to collect and are available at a much larger scale. Our key insight is to design a compatible spectral BRDF representation, along with a training strategy that leverages large-scale RGB BRDF data to facilitate spectral BRDF generation.
@@
In this paper, we propose SpecGen, which, to the best of our knowledge, is the first network to generate spectral BRDFs from a single RGB image. The core of SpecGen is the Spectral-Spatial Tri-plane Aggregation (SSTA) module, which facilitates the mapping between RGB and spectral BRDFs. Specifically, spectral BRDFs can be represented as a 4D function of $({\theta}_{h},{\theta}_{d},{\varphi}_{d},\lambda)$, where the spatial attributes $({\theta}_{h},{\theta}_{d},{\varphi}_{d})$—related to incident and outgoing directions—are shared between spectral and RGB BRDFs, while the spectral attribute $\lambda$ (wavelength) is unique to spectral BRDFs. Inspired by the 4D function decomposition in dynamic neural radiance fields $(x,y,z,t)$, we decompose the spatial and spectral attributes into two tri-planes, encoding their corresponding latent features separately. Building upon this tri-plane representation, we design a training strategy that leverages large-scale RGB BRDF datasets to supervise the spatial tri-planes, thereby alleviating the data scarcity issue in the spectral domain. Furthermore, to effectively integrate RGB and spectral features, we propose an Adaptive Feature Fusion (AFF) module, which replaces traditional feature fusion methods such as the Hadamard product [15]. The AFF module enhances the model’s representational capability by adaptively combining multi-plane features. As shown in Fig. 1, these dedicated designs enable our method to accurately generate spectral BRDFs and to facilitate downstream spectral rendering tasks.
To summarize, our main contributions are as follows:
@@
II Related work
As shown in Table I, we summarize existing methods for BRDF generation and spectral image reconstruction from RGB images. Since our method is based on neural implicit representations, we also briefly review relevant works on spectral BRDF generation and neural implicit representations.
II-A BRDF Generation
BRDF reconstruction and generation play a key role in advancing computer graphics and material analysis. Early methods, such as those based on CNNs [16, 17], extract spatially varying BRDFs (SVBRDFs) from images, but struggle with generalization and require separate training for each material. Later works utilize U-Net architectures [18, 19], which improve generalization performance but often fail to preserve fine details, particularly in highlight regions. To address this limitation, GAN-based methods [20, 21, 22] have gained popularity. For instance, SurfaceNet [20] and MaterialGAN [21] utilize GAN structures to generate realistic SVBRDFs from a single image. TileGen [22] extends MaterialGAN by training category-specific models, while PhotoMat [23] builds upon prior works by incorporating real-world datasets to further enhance visual quality.
With the rise of diffusion models [24], recent work has explored diffusion-based approaches for BRDF generation, such as ControlMat [8] and MatFuse [9]. These methods address the low-quality artifacts observed in GAN-based approaches, but suffer from slower generation speeds. To improve inference efficiency, meta-learning-based methods [25, 26] have been proposed, achieving faster generation at the cost of reduced accuracy. Whereas all aforementioned methods focus solely on RGB BRDF generation, our approach, to the best of our knowledge, is the first to enable both RGB and spectral BRDF representation and generation.
The above BRDF generation methods are based on parametric reflectance models [27], representing reflectance via parameters such as albedo, roughness, and metallicity. Another line of research focuses on measured BRDFs, such as the MERL [14] and RGL [6] datasets. Matusik et al. [14] propose a data-driven BRDF generation model using PCA basis extraction. Gokbudak et al. [7] introduce a continuous neural BRDF representation for sparse reconstruction of BRDFs for unseen materials. Narumoto et al. [28] propose a neural embedding for BRDFs and a temporal model to represent BRDF changes in latent space. However, these methods primarily focus on RGB BRDFs, leaving the problem of spectral BRDF generation largely unexplored.
Table I: Summary of existing representative methods on BRDF generation.
@@ II-A BRDF Generation
ControlMat [8], MatFuse [9] RGB Image RGB SVBRDF Diffusion Model
PhotoMat [23], TileGen [22] Flash RGB Image RGB SVBRDF GAN
Metappearance [25] Image & Light RGB BRDF Meta Learning
Gokbudak et al. [7] BRDF Sampling Data RGB BRDF Hyper Network
Narumoto et al. [28] BRDF Sampling Data Time-varying RGB BRDF Coordinate-based MLP
@@
II-B Implicit Neural Representations
@@
Unlike traditional discrete representations—such as meshes, voxels, or point grids—which store sampled values at fixed locations, incurring substantial memory costs, suffering from aliasing, and requiring explicit interpolation schemes to achieve smoothness [29]—implicit neural representations (INRs) employ a learnable function (typically an MLP) that directly maps each continuous coordinate in the signal domain to its corresponding signal value. This formulation endows INRs with several attractive properties: they are resolution-independent, differentiable, and drastically more memory-efficient than high-resolution grids. A growing body of work validates these advantages across diverse domains. For example, SIREN [30] uses sinusoidal activations to capture high-frequency natural signals and their derivatives; NeRF [31] represents static scenes as continuous 5-D radiance fields for photorealistic novel-view synthesis; Neural Descriptor Fields (NDFs) [32] encode object points along with their relative poses, enabling smooth interpolation of robot grasp trajectories; and Local Implicit Neural Representation (LINR) [33] modules enhance fine-grained visual tracking by refining target regions with coordinate-based implicit functions.
@@
Building on this foundation, K-Planes [15] replaces the monolithic MLP with a set of locally parameterized feature planes, thereby reducing the memory requirement from $O(N^{4})$ to $O(6N^{2})$ for four-dimensional signals and enabling asymmetric supervision across coordinate subsets. Inspired by K-Planes, our SpecGen introduces a novel spectral and spatial triplane representation to effectively decompose and model spectral BRDFs, thereby addressing the challenge of data scarcity in spectral BRDF generation.
II-C Hyperspectral Image Reconstruction
@@
Hyperspectral image (HSI) reconstruction aims to recover a full spectral data cube—comprising two spatial dimensions and one spectral dimension—from inputs with limited spectral resolution, such as RGB photos. Li et al. [10] show that deep CNNs can enhance textures and synthesize plausible spectra. Wang et al. [11] further improve robustness by framing the task as an optimization problem and adopting data-driven priors. More recently, Cai et al. [34] employ a mask-guided spectral transformer with self-attention to capture long-range spatial–spectral relationships, achieving state-of-the-art results.
@@
Despite these advances, existing HSI methods operate strictly in the image domain: they “uplift” the per-pixel spectral signatures of the input frame but cannot disentangle the underlying material reflectance from lighting and geometry. Consequently, the recovered spectra cannot be transferred to new shapes, viewing directions, or illumination environments, limiting their usefulness in downstream graphics and vision tasks such as relighting, compositing, or physically based simulation. In contrast, our spectral-BRDF generation paradigm explicitly models the bidirectional and wavelength-dependent reflectance of the material itself, enabling faithful spectral rendering on arbitrary geometries and under arbitrary lighting.
III Method
@@
As shown in Fig. 2, given an input RGB image, the SSTA network leverages an encoder-decoder to extract two sets of triplanes: a spatial triplane that encodes the reflectance response of incident and outgoing light angles, and a spectral triplane that captures the reflectance responses jointly across angles and wavelengths. Next, we project wavelengths and incident-outgoing angles, represented in Rusinkiewicz coordinates [35], onto the two triplanes to generate six feature vectors. The AFF module then fuses these features and produces a latent BRDF feature, which is used as input for our MLP-based BRDF mapping module to predict the corresponding spectral reflectance $r$.
@@
Figure 2: Pipeline of our method.
III-A Spectral-Spatial Tri-plane Aggregation
Insight: K-Planes [15] for Dynamic Neural Radiance Field.
Dynamic neural radiance fields can be represented as a 4D function with respect to 3D spatial location $(x,y,z)$ and a 1D timestamp $t$. Therefore, K-Planes [15], as a dynamic NeRF method, models the dynamic scene radiance $c$ as
@@ III-A Spectral-Spatial Tri-plane Aggregation
To decompose the static canonical scene and dynamic motion field, K-Planes [15] introduces a planar factorization that represents this 4D function using $C_{4}^{2}$ feature planes. Among these, three planes $\left\{(x,y),\ (x,z),\ (y,z)\right\}$ encode spatial domain information, while the remaining three $\left\{(x,t),\ (y,t),\ (z,t)\right\}$ capture spatial-temporal variations, thus enabling a clear decoupling of the static radiance field and temporal motion field.
@@
Based on Rusinkiewicz coordinates [35], the spectral BRDF response can also be modeled as a 4D function of the incident-outgoing angles $(\theta_{h},\theta_{d},\varphi_{d})$ and wavelength $\lambda$. Specifically, given incident and outgoing light directions $\boldsymbol{\omega_{i}}$ and $\boldsymbol{\omega_{o}}$, the half vector can be calculated as $\boldsymbol{h}=\frac{\boldsymbol{\omega_{i}}+\boldsymbol{\omega_{o}}}{\|\boldsymbol{\omega_{i}}+\boldsymbol{\omega_{o}}\|}$. The angle $\theta_{h}$ denotes the polar angle of $\boldsymbol{h}$ relative to the fixed surface normal $\boldsymbol{n}=(0,0,1)$; $\theta_{d}$ is the polar angle between the outgoing direction $\boldsymbol{\omega_{o}}$ and the half-vector $\boldsymbol{h}$; $\varphi_{d}$ is the azimuth angle around $\boldsymbol{h}$. Similar to the 4D function in Dynamic NeRF, the spectral BRDF response $r$ can be denoted as:
@@
As shown in Fig. 2, we follow the decomposition in K-Planes [15] to decompose the 4D spectral response into two triplanes. The three planes $\{(\theta_{h},\theta_{d}),(\theta_{h},\varphi_{d}),(\theta_{d},\varphi_{d})\}$, which depend solely on angular parameters, are designated as the spatial triplane. We denote these three planes as $\boldsymbol{F}_{d1}$, $\boldsymbol{F}_{d2}$, and $\boldsymbol{F}_{d3}$, respectively. Meanwhile, the remaining three planes $\{(\theta_{h},\lambda),(\theta_{d},\lambda),(\varphi_{d},\lambda)\}$ incorporate wavelength dependence and are referred to as the spectral triplane, denoted by $\boldsymbol{F}_{e1}$, $\boldsymbol{F}_{e2}$, and $\boldsymbol{F}_{e3}$. Since the spatial triplane is only related to incident-outgoing angles, it can be shared by both RGB and spectral BRDFs. As discussed in our RGB-spectral joint training strategy, this decomposed representation enables the use of RGB data to enhance spectral BRDF generation.
@@
III-B Adaptive Feature Fusion (AFF)
Unlike previous methods [15, 36] that conduct feature fusion via dot product, we propose an Adaptive Feature Fusion (AFF) module, inspired by SKNet [37], to dynamically select appropriate weights for fusing different features. As shown in Fig. 2, we first reshape the $C$-dimensional 1D feature vectors into 3D feature maps of size $1\times H\times W$. Then, we apply $3\times 3$ convolutions and element-wise addition to promote feature interaction, resulting in fused hybrid features, denoted as $\boldsymbol{f}_{h}\in\mathbb{R}^{1\times H\times W}$:

### v2

@@ Abstract
Synthesizing spectral images across different wavelengths is essential for photorealistic rendering. Unlike conventional spectral uplifting methods that convert RGB images into spectral ones, we introduce SpecGen, a method that generates spectral bidirectional reflectance distribution functions (BRDFs) from a single RGB image of a sphere. This enables spectral image rendering under arbitrary illuminations and shapes covered by the corresponding material. A key challenge in spectral BRDF generation is the scarcity of measured spectral BRDF data. To address this, we propose the Spectral-Spatial Tri-plane Aggregation (SSTA) network, which models reflectance responses across wavelengths and incident-outgoing directions, allowing the training strategy to leverage abundant RGB BRDF data to enhance spectral BRDF generation. Experiments show that our method accurately reconstructs spectral BRDFs from limited spectral data and surpasses existing methods in hyperspectral image reconstruction, achieving an improvement of 8 dB in PSNR. Codes link: https://github.com/sosjzy/SpecGen.
1 Introduction
Real light is composed of a spectrum of wavelengths, and photorealistic image synthesis requires capturing the wavelength-dependent reflectance properties of materials under varying illumination and viewing conditions. Spectral bidirectional reflectance distribution functions (BRDFs) constitute a crucial ingredient in this process, enabling the synthesis of hyperspectral images for remote sensing [ghassemian2016review, cheng2022multi], material analysis [choudhary2022recent], and virtual reality [kang2020deephandsvr]. However, acquiring high-quality spectral BRDF data is challenging, as the measurement process requires scanning a four-dimensional domain at high resolution—a tedious and time-consuming task [dupuy2018adaptive], leading existing spectral BRDF datasets to be scarce and limited in scale. To bypass the need for extensive spectral measurements, it is desired to develop a spectral BRDF generation method that can produce photorealistic spectral BRDFs from a single image input.
@@ 1 Introduction
Figure 1: From a single RGB image of a sphere (left), our method recovers the spectral BRDF of the corresponding material, allowing the spectral image rendering of arbitrary illuminations and shapes covered by the material under different wavelengths (right).
Most existing BRDF generation methods target RGB BRDFs. For instance, Gokbudak et al[gokbudak2025hypernetworks] use a hypernetwork for sparse RGB BRDF reconstruction, while ControlMat [vecchio2024controlmat] and MatFuse [vecchio2024matfuse] employ diffusion models to predict spatially varying BRDF maps from a single flash RGB photo. These approaches, though effective for RGB, cannot be directly extended to spectral BRDFs due to representation differences and scarce spectral data. Separately, spectral uplifting methods [li2017hyperspectral, Wang_2019_CVPR, liu2024physics, chen2023spectral] estimate spectral images from RGB, but do not recover spectral BRDFs, hindering appearance transfer across shapes and lighting. Thus, despite advances in RGB BRDF and spectral image generation, spectral BRDF generation remains unsolved.
@@
The key challenge of spectral BRDF generation lies in the scarcity of spectral BRDF data. Currently, only 51 isotropic spectral BRDFs are publicly available [dupuy2018adaptive], which makes it difficult to train high-quality, generalizable models. In contrast, measured RGB BRDF datasets, such as MERL [matusik2003data], are relatively easy to collect and are available at a much larger scale. Our key insight is to design a compatible spectral BRDF representation, along with a training strategy that leverages large-scale RGB BRDF data to facilitate spectral BRDF generation.
@@
In this paper, we propose SpecGen, which, to the best of our knowledge, is the first network to generate spectral BRDFs from a single RGB image. The core of SpecGen is the Spectral-Spatial Tri-plane Aggregation (SSTA) module, which facilitates the mapping between RGB and spectral BRDFs. Specifically, spectral BRDFs can be represented as a 4D function of $({\theta}_{h},{\theta}_{d},{\varphi}_{d},\lambda)$, where spatial attributes $({\theta}_{h},{\theta}_{d},{\varphi}_{d})$—related to incident and outgoing directions—are shared between spectral and RGB BRDFs, while the spectral attribute $\lambda$ (wavelength) is unique to spectral BRDFs. Inspired by the 4D function decomposition in dynamic neural radiance fields $(x,y,z,t)$, we decompose spatial and spectral attributes into two tri-planes, encoding their corresponding latent features separately. Building upon this tri-plane representation, we design a training strategy that leverages large-scale RGB BRDF datasets to supervise the spatial tri-planes, thereby alleviating the data scarcity issue in the spectral domain. Furthermore, to effectively integrate RGB and spectral features, we propose an Adaptive Feature Fusion (AFF) module, which replaces traditional feature fusion such as Hadamard product [fridovich2023k]. The AFF module enhances the model’s representation by adaptively combining multi-plane features. As shown in Fig. 1, these dedicated designs enable our method to accurately generate spectral BRDFs and to facilitate downstream spectral rendering tasks. Our main contributions are as follows:
@@
2 Related work
As shown in Table 1, we summarize existing methods for BRDF generation and spectral image reconstruction from RGB images. Since our method is based on neural implicit representations, we briefly review relevant works on spectral BRDF generation and neural implicit representations.
Table 1: Summary of existing representative methods on BRDF generation.
@@ 2 Related work
ControlMat [vecchio2024controlmat], MatFuse [vecchio2024matfuse] RGB Image RGB SVBRDF Diffusion Model
PhotoMat [zhou2023photomat], TileGen [zhou2022tilegen] Flash RGB Image RGB SVBRDF GAN
Metappearance [fischer2022metappearance] Image & Light RGB BRDF Meta Learning
Gokbudak et al[gokbudak2025hypernetworks] BRDF Sampling Data RGB BRDF Hyper Network
Narumoto et al[narumoto2024synthesizing] BRDF Sampling Data Time-varying RGB BRDF Coordinate-based MLP
@@
2.1 BRDF Generation
BRDF reconstruction and generation play a key role in advancing computer graphics and material analysis. Early methods, such as those based on CNNs [deschaintre2020guided, li2017modeling], extract spatially varying BRDFs (SVBRDFs) from images, but struggle with generalization and require separate training for each material. Later works utilize U-Net architectures [deschaintre2019flexible, deschaintre2018single], which improve generalization performance but often fail to preserve fine details, particularly in highlight regions. To address this limitation, GAN-based methods [vecchio2021surfacenet, guo2020materialgan, zhou2022tilegen] have gained popularity. For instance, SurfaceNet [vecchio2021surfacenet] and MaterialGAN [guo2020materialgan] utilize GAN structures to generate realistic SVBRDFs from a single image. TileGen [zhou2022tilegen] extends MaterialGAN by training category-specific models, while PhotoMat [zhou2023photomat] builds upon prior works by incorporating real-world datasets to further enhance visual quality.
With the rise of diffusion models [ho2020denoising], recent works such as ControlMat [vecchio2024controlmat] and MatFuse [vecchio2024matfuse] have applied them to BRDF generation, effectively reducing the low-quality artifacts common in GAN-based approaches but at the expense of markedly slower generation speeds. To improve efficiency, meta-learning methods [fischer2022metappearance, zhou2022look] enable rapid inference within limited computational budgets, though this often comes with a noticeable drop in accuracy. Despite these advances, all above approaches focus exclusively on RGB BRDFs; to our knowledge, ours is the first to enable both RGB and spectral BRDF representation and generation.
Most existing BRDF generation approaches rely on parametric reflectance models [burley2012physically], parameterizing reflectance via attributes such as albedo, roughness, and metallicity. Another line of work explores measured BRDFs from datasets like MERL [matusik2003data] and RGL [dupuy2018adaptive]. For example, Matusik et al[matusik2003data] used PCA to build a data-driven BRDF model; Gokbudak et al[gokbudak2025hypernetworks] proposed a continuous neural representation for sparse reconstruction of unseen materials; and Narumoto et al[narumoto2024synthesizing] developed a neural embedding with a temporal model to capture BRDF changes in latent space. However, these also concentrate on RGB BRDFs, leaving spectral BRDF generation largely unexplored.
2.2 Implicit Neural Representations
@@
Unlike discrete representations such as meshes, voxels, or point grids—which store sampled values at fixed locations, leading to high memory usage, aliasing, and the need for explicit interpolation [hanocka2019meshcnn]—implicit neural representations (INRs) use a learnable function (typically an MLP) to map each continuous coordinate in the signal domain to its value. This design makes INRs resolution-independent, differentiable, and far more memory-efficient than dense grids. These advantages have been demonstrated in various domains: SIREN [sitzmann2020implicit] employs sinusoidal activations to capture high-frequency signals and derivatives; NeRF [mildenhall2021nerf] models static scenes as continuous 5-D radiance fields for photorealistic novel-view synthesis; and Neural Descriptor Fields (NDFs) [simeonov2022neural] encode object points with relative poses for smooth interpolation of robot grasp trajectories.
@@
Building on this foundation, K-Planes [fridovich2023k] replaces the monolithic MLP with a set of locally parameterized feature planes, thereby reducing the memory requirement from $O(N^{4})$ to $O(6N^{2})$ for four-dimensional signals and enabling asymmetric supervision across coordinate subsets. Inspired by K-Planes, our SpecGen introduces a novel spectral and spatial triplane representation to effectively decompose and model spectral BRDFs, thereby addressing the challenge of data scarcity in spectral BRDF generation.
2.3 Hyperspectral Image Reconstruction
@@
Hyperspectral image (HSI) reconstruction seeks to recover a full spectral data cube—from inputs with limited spectral resolution such as RGB images. Li et al[li2017hyperspectral] showed that deep CNNs can enhance textures and synthesize plausible spectra; Wang et al[Wang_2019_CVPR] improved robustness via optimization with data-driven priors; and Cai et al[cai2022mask] used a mask-guided spectral transformer with self-attention to capture long-range spatial–spectral relationships, achieving the best results.
@@
However, existing HSI methods [li2017hyperspectral, Wang_2019_CVPR, cai2022mask] operate solely in the image domain, uplifting per-pixel spectra without disentangling material reflectance from lighting and geometry, preventing transfer to new shapes, viewpoints, or illumination. This limits applications such as relighting, compositing, and physically based simulation. In contrast, our spectral-BRDF paradigm explicitly models bidirectional, wavelength-dependent material reflectance, enabling accurate spectral rendering on arbitrary geometries and arbitrary lighting.
3 Method
@@
As shown in Fig. 2, given an input RGB image, the SSTA network leverages an encoder-decoder to extract two sets of triplanes: a spatial triplane that encodes the reflectance response of incident and outgoing light angles, and a spectral triplane that captures the reflectance responses jointly across angles and wavelengths. Next, we project wavelengths and incident-outgoing angles, represented in Rusinkiewicz coordinates [rusinkiewicz1998new], onto the two triplanes to generate six feature vectors. The AFF module then fuses these features and produces a latent BRDF feature, which is used as input for our MLP-based BRDF mapping module to predict the corresponding spectral reflectance $r$.
@@
Figure 2: Pipeline of our SpecGen.
3.1 Spectral-Spatial Tri-plane Aggregation
Insight: K-Planes [fridovich2023k] for Dynamic Neural Radiance Field.
Dynamic neural radiance fields can be represented as a 4D function with respect to 3D spatial location $(x,y,z)$ and a 1D timestamp $t$. Therefore, K-Planes [fridovich2023k], as a dynamic NeRF, models the dynamic scene radiance $c$ as
@@ 3.1 Spectral-Spatial Tri-plane Aggregation
To decompose the static canonical scene and dynamic motion field, K-Planes [fridovich2023k] introduces a planar factorization that represents this 4D function using $C_{4}^{2}$ feature planes. Among these, three planes $\left\{(x,y),\ (x,z),\ (y,z)\right\}$ encode spatial domain information, while the remaining three $\left\{(x,t),\ (y,t),\ (z,t)\right\}$ capture spatial-temporal variations, thus enabling a clear decoupling of the static radiance field and temporal motion field.
@@
Based on Rusinkiewicz coordinates [rusinkiewicz1998new], the spectral BRDF response can also be modeled as a 4D function of the incident-outgoing angles $(\theta_{h},\theta_{d},\varphi_{d})$ and wavelength $\lambda$. Specifically, given incident and outgoing light directions $\boldsymbol{\omega_{i}}$ and $\boldsymbol{\omega_{o}}$, the half vector can be calculated as $\boldsymbol{h}=\frac{\boldsymbol{\omega_{i}}+\boldsymbol{\omega_{o}}}{\|\boldsymbol{\omega_{i}}+\boldsymbol{\omega_{o}}\|}$. The angle $\theta_{h}$ denotes the polar angle of $\boldsymbol{h}$ relative to the fixed surface normal $\boldsymbol{n}=(0,0,1)$; $\theta_{d}$ is the polar angle between the outgoing direction $\boldsymbol{\omega_{o}}$ and the half-vector $\boldsymbol{h}$; $\varphi_{d}$ is the azimuth angle around $\boldsymbol{h}$. Similar to the 4D function in Dynamic NeRF, the spectral BRDF response $r$ can be denoted as:
@@
As shown in Fig. 2, we follow the decomposition in K-Planes [fridovich2023k] to decompose the 4D spectral response into two triplanes. The three planes $\{(\theta_{h},\theta_{d}),(\theta_{h},\varphi_{d}),(\theta_{d},\varphi_{d})\}$, which depend solely on angular parameters, are designated as the spatial triplane. We denote these three planes as $\boldsymbol{F}_{d1}$, $\boldsymbol{F}_{d2}$, and $\boldsymbol{F}_{d3}$, respectively. Meanwhile, the remaining three planes $\{(\theta_{h},\lambda),(\theta_{d},\lambda),(\varphi_{d},\lambda)\}$ incorporate wavelength dependence and are referred to as the spectral triplane, denoted by $\boldsymbol{F}_{e1}$, $\boldsymbol{F}_{e2}$, and $\boldsymbol{F}_{e3}$. Since the spatial triplane is only related to incident-outgoing angles, it can be shared by both RGB and spectral BRDFs. As discussed in our RGB-spectral joint training, this decomposed representation enables the use of RGB data to enhance spectral BRDF generation.
@@
3.2 Adaptive Feature Fusion (AFF)
Unlike previous methods [fridovich2023k, li2025dual] that conduct feature fusion via dot product, we propose an Adaptive Feature Fusion (AFF) module, inspired by SKNet [li2019selective], to dynamically select appropriate weights for fusing different features. As shown in Fig. 2, we first reshape the $C$-dimensional 1D feature vectors into 3D feature maps of size $1\times H\times W$. Then, we apply $3\times 3$ convolutions and element-wise addition to promote feature interaction, resulting in fused hybrid features, denoted as $\boldsymbol{f}_{h}\in\mathbb{R}^{1\times H\times W}$: